By using Ohm’s Law, the current flowing through a resistor can be calculated as follows:

Calculate the current flowing through a 100Ω resistor that has one of its terminals connected to 50 volts and the other terminal connected to 30 volts.

Grade:11

1 Answers

Vikas TU
4 years ago
Dear student 
Calculate the current flowing through a 100Ω resistor that has one of its terminals connected to 50 volts and the other terminal connected to 30 volts.
Voltage at terminal A is equal to 50v and the voltage at terminal B is equal to 30v. Therefore, the voltage across the resistor is given as:
VA = 50v, VB = 30v, therefore, VA – VB = 50 – 30 = 20v
The voltage across the resistor is 20v, then the current flowing through the resistor is given as:
I = VAB ÷ R = 20V ÷ 100Ω = 200mA